OpenStudy (anonymous):

If the average of three numbers is V. If one of the numbers is Z and another is Y, what is the remaining number? A. ZY - V B. Z/V - 3 - Y C. Z/3 - V - Y D. 3V- Z - Y E. V- Z - Y

OpenStudy (anonymous):

How about it, axes18?

OpenStudy (anonymous):

D.3V-Z-Y

OpenStudy (radar):

D

OpenStudy (anonymous):

Explanation?

OpenStudy (radar):

3V=z+y+x x=3v-y-z

OpenStudy (anonymous):

Let the number be X. (X+Y+Z)/3 = V [Average of three] X+Y+Z = 3V X= 3V-Z-Y [ANS}

OpenStudy (anonymous):

Happy with that, tham?